Africa Traders is not a registered VAT vendor and buys and sells furniture. Africa Traders sold a dining room table for R3 400. The mark-up percentage on cost is 25%

1
Expert's answer
2021-09-22T09:24:44-0400

Sales = Cost + mark-up

When calculating cost and sale price, VAT is normally excluded.

Therefore, we start by first eliminating VAT from the sales:

Sales inclusive of 15"\%" VAT = 1,610


Sales exclusive of 15"\%" VAT = "\frac{100}{115} \times 1,610 = 1,400"


Now, determine cost:

Cost + mark-up = Sales

100"\%" + 20"\%" = 120"\%"

100"\%" + 20"\%" = 1,400


Therefore:

Cost of sales = "\frac{100}{120} \times 1,400 = 1,166.67 = 1,167"


The cost of sales for this transaction = R1,167